The employees of Xitrex, Inc., are paid each Friday. The companys fiscal year-end is June 30, which falls on a Wednesday forComplete this question by entering your answers in the tabs below. Required 1 Required 2 Prepare an adjusting entry to recordView transaction list Journal entry worksheet < 1213 Record the reversing entry on July 1. Note: Enter debits before credits.View transaction list Journal entry worksheet Record the payment of salaries on July 2. Note: Enter debits before credits. GeView transaction list Journal entry worksheet Record the accrued salaries as of June 30. Note: Enter debits before credits. GView transaction list Journal entry worksheet Record the payment of salaries on July 2 assuming a reversing entry is not madeView transaction list 1 Record the accrued salaries as of June 30. 2 Record the payment of salaries on July 2 assuming a reve
Answer #1
1
Date General Journal Debit Credit
June 30 Salaries and wages expense 13200 =22000*3/5
      Salaries and wages payable 13200
July 1 Salaries and wages payable 13200
   Salaries and wages expense 13200
July 2 Salaries and wages expense 22000
      Cash 22000
2
Date General Journal Debit Credit
June 30 Salaries and wages expense 13200 =22000*3/5
      Salaries and wages payable 13200
July 2 Salaries and wages expense 8800 =22000*2/5
Salaries and wages payable 13200
      Cash 22000